OpenStudy (anonymous):

I need to switch y=5x+11 from slop intercept to standard. How do you do that?I know the formula is Ax+By=C but what do A,B&C mean and how do i get them out of the slope intercept equation?

OpenStudy (anonymous):

*slope intercept

OpenStudy (anonymous):

So, y=5x+11 is y=mx+b, or in other words, slope-intercept form. If you want to change slope intercept form into standard form(Ax+By=C) then follow these steps. First get your stating point set up. \[y=mx+b\] \[y=5x+11\] now we need to stat working the problem, start by eliminating 11. \[y-11=5x+11-11\] \[y-11=5x\] Now we need to eliminate y, remember, when you do something to one side of an equation, you have to do the same thing to the other side. \[y-11-y=5x-y\] \[-11=5x-y\] This is the final answer, but it is not in the correct orientation, so place -11 on the opposite side of the equation. \[5x-y=-11\] There you go, please mark this as best response if it helped;)

OpenStudy (anonymous):

@mackenzie3160